OpenStudy (andu1854):

The line that is normal to the curve x^2 + 2xy -3y^2 = 0 at (6,6) intersects the curve at what other point. Using the derivatives of this I was able to determine that the normal line is y = -x+12. I am not sure what to do next to figure out the other point of the equation.

ganeshie8 (ganeshie8):

you want to find where this normal line intersects the curve again

OpenStudy (andu1854):

yes

ganeshie8 (ganeshie8):

one dumb way is to solve the system : x^2 + 2xy -3y^2 = 0 y = -x+12

ganeshie8 (ganeshie8):

plugin y = -x+2 into the first equation you will get a quadratic in x which you can solve by factoring

OpenStudy (andu1854):

Ok that makes sense and by doing this it will give me two sets of coordinates (6,6) and whatever the other point might be?

ganeshie8 (ganeshie8):

or if you're clever ennough you will see that the given curve is just a pair of intersecting lines and canc be factored into form : \[(y-m_1x)(y-m_2x)=0\]

ganeshie8 (ganeshie8):

you're right!

OpenStudy (andu1854):

Cool, thank you very much, sometimes I just need a little help with being pointed in the right direction...

ganeshie8 (ganeshie8):

you're welcome :)

OpenStudy (anonymous):

Okay, I'm surprised the straight forward approach wasn't pointed out here...

OpenStudy (anonymous):

Differentiate with respect to \(x\), then set \((x,y)=(6,6)\) and solve for \(y'\). This gives the tangent slope. The normal line is perpendicular, so it's slope will be \(-1/y'\).

ganeshie8 (ganeshie8):

the op already has a normal @wio